OpenStudy (anonymous):

What number must multiply each side of the equation 3/5x=12 to produce the equivalent equation x=20

OpenStudy (anonymous):

3/5 is a fraction just so u know

OpenStudy (anonymous):

/5)x = 12 (5/3)(3/5)x = (5/3)12 x = 20 Therefore, 5/3

OpenStudy (anonymous):

THANKS!!~!

OpenStudy (anonymous):

Solve. 1/3x=12

OpenStudy (anonymous):

plz 1 more